### v1

# Multi-Prompting Decoder Helps Better Language Understanding

Abstract
Recent Pre-trained Language Models (PLMs) usually only provide users with the inference APIs, namely the emerging Model-as-a-Service (MaaS) setting. To adapt MaaS PLMs to downstream tasks without accessing their parameters and gradients, some existing methods focus on the output-side adaptation of PLMs, viewing the PLM as an encoder and then optimizing a task-specific decoder for decoding the output hidden states and class scores of the PLM. Despite the effectiveness of these methods, they only use a single prompt to query PLMs for decoding, leading to a heavy reliance on the quality of the adopted prompt. In this paper, we propose a simple yet effective Multi-Prompting Decoder (MPD) framework for MaaS adaptation. The core idea is to query PLMs with multiple different prompts for each sample, thereby obtaining multiple output hidden states and class scores for subsequent decoding. Such multi-prompting decoding paradigm can simultaneously mitigate reliance on the quality of a single prompt, alleviate the issue of data scarcity under the few-shot setting, and provide richer knowledge extracted from PLMs. Specifically, we propose two decoding strategies: multi-prompting decoding with optimal transport for hidden states and calibrated decoding for class scores. Extensive experiments demonstrate that our method achieves new state-of-the-art results on multiple natural language understanding datasets under the few-shot setting. All the code and data of this work can be obtained at https://github.com/czlll/MPD.
Index Terms:
Model-as-a-Service, few-shot learning.
I Introduction
Pre-trained Language Models (PLMs) have recently achieved remarkable performance on various downstream language understanding tasks with the “pre-training and fine-tuning” paradigm [1]. However, as the scale of PLMs continues to grow, fine-tuning the full model for each downstream task has become computationally expensive and deployment-inefficient. In light of this, Model-as-a-Service (MaaS) [2, 3] has emerged as a more practical model deployment scheme, allowing pre-trained models to serve downstream tasks by providing inference APIs. Within the MaaS scheme, users can access the outputs of PLMs (e.g., hidden states, class scores, or textual output) through APIs [^1] Currently, some famous MaaS services include GPT-3.5 Turbo, which offers APIs for accessing class scores and text output, as well as text-embedding-3-small and text-embedding-3-large, which provide APIs for accessing embedding representations (i.e., hidden states) of input text., but cannot access the parameters and gradients of PLMs, presenting a challenge in effectively adapting PLMs to downstream tasks. As shown in Table I, MaaS belongs to the category of grey-box PLM adaptation methods, which, unlike white-box adaptation methods, don’t require access to model parameters.
TABLE I: Comparison of existing PLM adaptation methods. Compared to kNN-Adapter, MaaS focuses on downstream task adaptation and does not require retrieval corpus.
Methods w/o Model Parameters w/o Hidden State w/o Token Probabilities w/o Retrieval Corpus
   White-Box LLM Fine-Tuning: full access to PLM
Fine-Tuning
Adapter [4]
Prefix-Tuning [5]
LoRA [6]
   Grey-Box LLM Adaptation: can not access the parameters of PLM
kNN-Adapter [7]
MaaS [3]
Figure 1: Performance and variation using five different prompts on four datasets.
To adapt MaaS PLMs to downstream tasks, existing methods have primarily explored both the input-side and the output-side adaptation. On the input-side adaptation, researchers focus on utilizing gradient-free methods to heuristically search a good continuous [3, 8] or discrete prompt [2, 9] for the target task, such as using evolutionary algorithm and reinforcement learning. However, input-side methods need to query PLM thousands of times for optimization, resulting in significant time overhead and the optimization process is very difficult due to the huge search space and the lack of gradients. On the output-side adaptation, researchers [10, 11] treat the PLM as an encoder and locally train a task-specific decoder to post-process PLM’s outputs. Compared to the input-side adaptation, the output-side adaptation methods only query PLM a few times and can benefit from gradient for optimization, often resulting in better performance.
Despite the effectiveness of these output-side adaptation methods, they usually only adopt a single prompt to query PLMs for a certain downstream task, thus often struggling with the performance variations brought by using different prompts [12, 13]. As shown in Figure 1, on all four datasets, the maximum performance fluctuations in accuracy can exceed 8% when using different prompts. Therefore, the quality of prompt is crucial, and relying on only one prompt easily exposes them to performance risks of encountering a low-quality query.
To this end, we propose to query PLMs with multiple prompts for a certain downstream task. This multi-prompting paradigm offers several advantages. Firstly, using multiple prompts helps mitigate reliance on the quality of a single prompt, leading to better stability. Secondly, in the few-shot setting, using multiple prompts enables a single sample to obtain multiple representations, thereby addressing the issue of data scarcity. Lastly, multiple prompts can guide the extraction of knowledge from PLMs from different perspectives, providing richer knowledge for downstream tasks.
Based on the above analysis, in this paper, we propose a novel output-side PLMs adaptation framework for few-shot classification, named Multi-Prompting Decoder (MPD). This MPD framework can decode the outputs of PLMs, including hidden states and class scores, to yield classification results. In contrast to previous methods, MPD adopts a novel multi-prompting decoding paradigm, which queries PLMs with multiple prompts for a sample, thereby obtaining more representation information for decoding. Specifically, we propose two decoding strategies: multi-prompting decoding with optimal transport for hidden states and calibrated multi-prompting decoding for class scores. Extensive experiments demonstrate that our method achieves new state-of-the-art results under the few-shot setting on multiple natural language understanding tasks.
In summary, the major contributions of this paper are:
  • We propose an output-side PLMs adaptation framework, multi-prompting decoder, for few-shot classification.
  • We design two multi-prompting decoding strategies: multi-prompting decoding with optimal transport and calibrated multi-prompting decoding to decode hidden states and class scores respectively.
  • Extensive experiments demonstrate that our method achieves new state-of-the-art results under the few-shot setting on multiple natural language understanding tasks, including sentiment analysis, topic classification, and natural language inference. Furthermore, our framework is demonstrated to be lightweight, efficient, and applicable to PLMs of different architectures (e.g., encoder only, decoder only, and encoder-decoder).
The rest of the article is organized as follows. Section II provides a comprehensive review of related work on prompt tuning, MaaS adaptation, parameter-efficient tuning, and optimal transport. Section III describes the preliminaries about MaaS adaptation and optimal transport. Section IV describes the technical details of our proposed MPD. Experimental settings, comparison, and analyses are presented in Section V. Finally, we conclude our work in Section VI.
II Related Work
Our work focuses on how to efficiently adapt PLMs to downstream tasks. We review four lines of research: prompt tuning (data efficiency), MaaS adaptation methods, parameter-efficient tuning (parameter efficiency), and optimal transport respectively.
II-A Prompt Tuning
Prompt tuning aims to reduce the gap between PLM pre-training and fine-tuning via predicting [MASK] token in the prompt for data-efficiency and achieves great performance in zero-shot and few-shot learning [1]. Prompt-based tuning is originally applied in knowledge probing [14, 15], and it has been adopted broadly in NLP [16, 17, 18, 19]. Since prompt has a large impact on the performance of prompt tuning, many studies have focused on learning a good prompt to alleviate the prompt engineering, including discrete prompt [20, 21, 22] and continuous prompt [23, 24, 5, 25, 26]. However, optimizing prompts is difficult [25] and sometimes computationally expensive and not interpretable. Our work focuses on output-side adaptation and uses manual prompts to make PLM adaptable to downstream tasks.
II-B MaaS Adaptation
In MaaS settings, users can access the outputs of PLMs (e.g., hidden states, class scores, or textual output) via APIs, but cannot access the parameters and gradients of PLMs. Thus, some works try to adapt these models to downstream tasks without accessing the model parameters and the gradients [3, 2].
Some input-side adaptation work about MaaS uses reinforcement learning [2, 9] or derivative-free optimization [3, 8] to find the optimal prompt. For the first type of method, Diao et al. [2] first use a variance-reduced policy gradient algorithm to estimate the gradient of the prompt token distribution. Deng et al. [9] use a policy network and reward function to generate and optimize discrete prompts. TEMPERA [27] constructs query-dependent prompts through test-time prompt editing and formulates this as a reinforcement learning problem. DP2[28] first designs a prompt generation strategy through multi-round dialogue alignment on GPT-4 and prompt evaluation metric SUE. Then, DP2O constructs a reinforcement learning framework based on policy gradients to match the prompts to inputs optimally. For the second type of method, BBT [3] adopts a covariance matrix adaptation evolution algorithm to optimize continuous prompt. BBTv2 [8] further prepends continuous prompts to every layer of PLM and proposes a divide-and-conquer gradient-free algorithm to optimize them. GDFO [29] uses a prompt generator to generate an extra vector as input based on the BBT framework. However, searching in a large space of prompts is inefficient and difficult to optimize.
Inspired by this, output-side adaptation methods use manual prompts and train task-specific decoder to post-process PLM’s outputs. PromptBoosting [10] constructs a series of weak learners, learns the corresponding verbalizer, and uses AdaBoost algorithm to ensemble them. DecT [11] extracts text feature and optimizes class prototypes as hyperspheres with a radius parameter.
II-C Parameter-Efficient Tuning
Parameter-Efficient Tuning (PET) aims to adjust only a small number of parameters while keeping most pre-trained parameters frozen. Houlsby et al. [4] first add two lightweight adapter modules to every layer. Afterwards, prefix tuning [5] prepends multiple trainable prefix vectors to the keys and values of the multi-head attention at every layer and LoRA [6] injects trainable low-rank matrices into every layer to approximate the weight updates. Besides, Lester et al. [25] prepend continuous prompt tokens to the input layer, Zaken et al. [30] only optimize the bias vectors, and He et al. [31] present a unified framework that establishes connections between those PET methods. Despite their success with few parameter updates, they still require model gradients to update the small portion of parameters, which are unavailable in the MaaS setting.
II-D Optimal Transport
OT theory [32] was originally used to study the movement of many items from one place to another with minimal cost and provides a distance measure method between two distributions in a closed way. Due to the excellent properties of the distance measure, OT is used in many areas, such as topic model [33], legal case matching [34], vision-language models [35], and image classification [36]. In this paper, we use OT to align text features obtained using multiple prompts and multiple class prototypes.
III Preliminaries
We present two preliminaries that are relevant to our work, i.e., model-as-a-service adaptation and optimal transport.
III-A MaaS Adaptation
MaaS adaptation aims to adapt the PLMs to downstream tasks based on a given few-shot training set. Specifically, the setting needs to correctly predict the label $y\in\{1,\dots,N\}$ for test sample $x$ based on a given few-shot training set $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{NK}$ and PLM $\mathcal{M}$, where $N$ is the number of classes and each class has $K$ training samples. In the MaaS setting, we can only query the PLM model $\mathcal{M}$ with input $x$ and get corresponding outputs (e.g., hidden states and class scores) without accessing the model parameters and the gradients.
To better utilize the PLM, we use prompt learning to wrap input samples into prompts using templates. Specifically, we first use a template $\mathcal{T}$ with a [MASK] token to enclose each input. Then, we can query $\mathcal{M}$ with prompt $\mathcal{T}(x)$ to get the hidden states and class scores. Hidden states can be obtained by the representation of [MASK] token of the final layer and class scores $\mathbf{s}\in\mathbb{R}^{N}$ can be obtained by using the verbalizer to map predicted tokens at the [MASK] token to classes. Take sentiment analysis as an example, we can use
$\mathcal{T}(x)=x\ \text{In summary, it was {[MASK]}.}$] .
as prompt with verbalizer that maps “bad” and “great” as label words for negative and positive sentiment.
III-B Optimal Transport
OT aims to find the minimal cost transport plan between two distributions. In this paper, we focus on discrete distribution and define two discrete distributions as follows:
[…]
where $\bm{u}$ and $\bm{v}$ are the discrete probability vectors that sum to 1, and $\delta_{\bm{f}}$ is a Dirac delta function operating on $\bm{f}$ in the embedding space. Then, the cost matrix $\bm{C}\in\mathbb{R}^{P\times Q}$ in OT can be defined as $\bm{C}_{p,q}=1-\text{sim}(\bm{f}_{p},\bm{g}_{q})$ and each element in the cost matrix denotes the cost between $\bm{f}_{p}$ and $\bm{g}_{q}$, where sim(,) denotes a similarity metric function. The optimization problem of optimal transport is formulated as:
$\underset{\bm{T}}{\text{minimize}}\quad\sum_{p=1}^{P}\sum_{q=1}^{Q}\bm{T}_{p,q}\bm{C}_{p,q}$ (2)
$\text{subject to}\ \bm{T}\bm{1}_{Q}=\bm{u},\ \bm{T}^{\top}\bm{1}_{P}=\bm{v},\ \bm{T}\in\mathbb{R}^{P\times Q}_{+}.$
where $\bm{T}$ is the learned transport plan denotes the matching flow between two distributions and $\bm{T}_{p,q}$ is the amount of mass that needs to move $\bm{f}_{p}$ to $\bm{g}_{q}$.
Since direct optimization of the above processes is usually time-consuming, Cuturi et al. [37] use Sinkhorn distance which adds an entropic constraint for fast optimization. Thus, the new optimization objective is formulated as:
$\underset{\bm{T}}{\text{minimize}}\quad\sum_{p=1}^{P}\sum_{q=1}^{Q}\bm{T}_{p,q}\bm{C}_{p,q}-\lambda h(\bm{T})$ (3)
$\text{subject to}\ \bm{T}\bm{1}_{Q}=\bm{u},\ \bm{T}^{\top}\bm{1}_{P}=\bm{v},\ \bm{T}\in\mathbb{R}^{P\times Q}_{+},$
where $h(\cdot)$ is entropy and $\lambda\geq 0$ is a regularization hyper-parameter. Then we can have a fast optimization solution to get optimal transport plan $\bm{T}^{*}$ with a few iterations as:
$\bm{T}^{*}=\text{diag}(\bm{u}^{(t)})\exp(-\bm{C}/\lambda)\text{diag}(\bm{v}^{(t)}),$ (4)
where $t$ denotes the iteration and in each iteration $\bm{u}^{(t)}=\bm{u}/(\exp(-\bm{C}/\lambda)\bm{v}^{(t-1)})$ and $\bm{v}^{(t)}=\bm{v}/\left(\exp(-\bm{C}/\lambda)^{\top}\bm{u}^{(t)}\right)$, with the uniform initiation $\bm{u}$ and $\bm{v}$, and $\bm{v}^{(0)}=\bm{1}_{Q}$.
Figure 2: Illustration of the multi-prompting decoding with optimal transport. We first use multiple prompts to query PLM. Then, we use optimal transport to get the optimal transport plan between features and prototypes. Finally, we compute OT scores and optimize network.
IV Methodology
In this section, we first present the overview of multi-prompting decoder framework in Section IV-A. Then, we introduce the technical details of multi-prompting decoding with optimal transport and calibrated multi-prompting decoding in Sections IV-B and IV-C. Finally, we summary the multi-prompting decoder framework in Section IV-D.
IV-A Overview of Multi-Prompting Decoder Framework
We introduce a novel output-side MaaS adaptation framework, Multi-Prompting Decoder (MPD), for few-shot classification. This framework decodes the outputs of PLMs, including hidden states and class scores, to classify. In contrast to previous decoding frameworks that query PLMs only once for a sample, we propose multi-prompting decoding paradigm, which queries PLMs with multiple prompts for a sample, thereby obtaining more representation information for decoding. Specifically, for decoding the PLMs’ output hidden states, we design a multi-prompting decoding strategy with optimal transport to better align representations under multiple prompts and multiple prototypes. For decoding the PLMs’ output class scores, we additionally design a calibrated multi-prompting decoding strategy to calibrate class scores under multiple prompts. These two decoding strategies are jointly used in the final decoding process.
IV-B Multi-Prompting Decoding with Optimal Transport
In our multi-prompting paradigm, each sample has multiple types of hidden states corresponding to multiple prompts. Building a unified prototype for these hidden states neglects the characteristics of each prompt, leading to a crude solution. Building multiple prototypes specific to each type of hidden state is also a sub-optimal solution, which will lead to an ensemble of multiple prototypes, and these prototypes cannot benefit from each other during training.
To fully exploit the advantages of the multi-prompting decoding paradigm, we establish multiple prototypes for each class and design multi-prompting decoding with optimal transport to decode hidden states. Firstly, we establish multiple prototypes shared across multiple prompts for each class. In this way, multiple prompts can be used for joint learning, and multiple prototypes can capture the fine-grained differences among different prompts. Subsequently, as each sample has multiple representations and each class has multiple prototypes, we employ optimal transport to determine the optimal matching between text representations and prototypes for classification. Considering the limited number of training samples available in the few-shot setting, the decoder is lightweight, comprising only a linear layer and several learnable class prototypes.
IV-B1 Querying with Multiple Prompts
Given a text $x_{i}$, we first use a series of templates $\mathcal{T}_{1},\cdots,\mathcal{T}_{P}$ with a [MASK] token to enclose the text and then query $\mathcal{M}$ with template-wrapped inputs $\mathcal{T}_{1}(x_{i})$, $\cdots$, $\mathcal{T}_{P}(x_{i})$ to get a series of initial text representations $\mathbf{h}_{i,1},\cdots,\mathbf{h}_{i,P}$.
Specifically, we take hidden states of the final layer at the [MASK] position as the initial text representations extracted by PLM $\mathcal{M}$. Then, we use a linear layer parameterized by $\mathbf{W}$ to project the initial representations to get text representations for classification:
$\mathbf{v}=\mathbf{Wh}$ (5)
It is worth noting that the text representation matrix of text $x_{i}$ is $\mathbf{V}_{i}\in\mathbb{R}^{P\times d}$.
IV-B2 Features-Prototypes Matching via OT
Our MPD framework contains multiple text representations and class prototypes. Therefore, how to assign the relationship between representation and prototype is an essential issue for optimization. We employ OT to better align multiple text representations and multiple prototypes of each class.
OT aims to find the optimal transport plan $\bm{T}\in\mathbb{R}^{P\times Q}$ that represents the fine-grained matching flow between text representations and class prototypes, where $P$ denotes the number of prompts and $Q$ denotes the number of prototypes. The optimal transport plan assigns a higher weight to the pair of text representation and prototype that are close to each other. We use Sinkhorn’s algorithm [37] to optimize the transport plan. Specifically, we fix text representations and class prototypes to optimize the transport plan.
IV-B3 Classifying with OT Scores
After getting the optimal transport plan $\bm{T}$, the optimal transport score between text $x_{i}$ and class $k$ can be computed according to the optimal transport plan and similarity matrix. Specifically,
$\text{Dec}(x_{i},k)=\sum_{m=1}^{P}\sum_{n=1}^{Q}\bm{T}_{i,k}^{m,n}\text{sim}(\mathbf{V}_{i,m},\mathbf{R}_{k,n})$ (6)
where Dec(,) denotes the optimal transport score between text $x_{i}$ and class $k$, $\bm{T}_{i,k}^{m,n}$ denotes the ($m$,$n$)-th element in the optimal transport plan between text $x_{i}$ and class $k$, sim(,) denotes cosine similarity, $\mathbf{V}_{i,m}$ is the text representations of $\mathcal{T}_{m}(x_{i})$, and $\mathbf{R}_{k,n}$ is the $n$-th prototype representation of the $k$-th class.
Finally, we use the softmax function over optimal transport scores between text representations and all classes to get the predicted probability distribution, and use cross-entropy loss to optimize the parameters. Specifically,
$\mathcal{L}=\sum_{i=1}^{NK}\frac{-1}{NK}\text{log}\frac{\text{exp}(\text{Dec}(x_{i},y_{i}))}{\sum_{j=1}^{N}\text{exp}(\text{Dec}(x_{i},j))}$ (7)
Algorithm 1 The training flow of MPD
1:The few-shot training set and PLM $\mathcal{M}$.
2:The parameters of linear layer $\mathbf{W}$ and prototypes $\mathbf{R}$
[…]
4:Randomly initialize class prototypes $\bm{R}$
5:Feed few-shot training set into PLM $\mathcal{M}$ to get features $\mathbf{H}_{1},\cdots,\mathbf{H}_{NK}$
6:for iteration $=1,2,\dots,T$ do
7:     Sample a feature $\mathbf{H}_{i}$ and use linear layer to get the text representations $\bm{V}_{i}$
[…]
9:     for $k=1,2,\dots,N$ do
10:         Calculate the cost matrix $\bm{C}_{ik}=\bm{1}-\bm{V}_{i}\bm{R}_{k}^{\top}$
11:         for $t_{in}=1,2,\dots,T_{in}$ do
12:              $Z=\exp(-\bm{C}_{ik}/\lambda)$
13:              $\bm{u}^{(t_{in})}=\bm{u}/(Z\bm{v}^{(t_{in}-1)})$
14:              $\bm{v}^{(t_{in})}=\bm{v}/(Z^{\top}\bm{u}^{(t_{in})})$
15:              $\Delta_{v}=\sum|\bm{v}^{(t_{in})}-\bm{v}^{(t_{in}-1)}|/Q$
16:              if $\Delta_{v}<\delta$ then
17:                  break
18:              end if
19:         end for
20:         $\bm{T}^{*}_{ik}=\text{diag}(\bm{u}^{(t)})\exp(-\bm{C}_{ik}/\lambda)\text{diag}(\bm{v}^{(t)})$
21:         Calculate the OT scores by Eq. (6)
22:     end for
[…]
24:     Update the parameters with cross-entropy loss by Eq. (7)
25:end for
26:return The well-trained linear layer $\mathbf{W}$ and prototypes $\mathbf{R}$
IV-C Calibrated Multi-Prompting Decoding
We also decode the class scores from PLM’s outputs as supplementary for the hidden states decoding to provide additional prior knowledge of the PLMs. As shown in Figure 3, we first expand the set of label words for the verbalizer, then calibrate the class scores for each prompt separately, and average the calibrated class scores of all prompts. Finally, we combine OT scores and calibrated class scores for joint decoding.
IV-C1 Label Words Expansion
Considering that many label words express the same class, such as ‘good’ and ‘great’ both conveying positive sentiment, we first use a simple and effective way to expand the label words.
Given the prediction layer of masked language model $\mathbf{W}_{MLM}\in\mathbb{R}^{|\mathcal{V}|*d}$, we can treat each item as a representation of a word and expand the label words based on that representation. Then, given a label word, we select the top-k words with the highest cosine similarity to it as the set of expanded label words S. Due to the possibility of noise in the set of expanded label words, we use the softmax function over cosine similarity to assign a weight to each element in the set to mitigate the noise. It is worth noting that due to the large number of open-sourced PLMs, getting an expanded set of label words is easy.
Figure 3: Illustration of the calibrated multi-prompting decoding.
IV-C2 Calibrating Class Scores
Predictions of PLM are usually biased because PLMs tend to predict tokens that are common in its pre-training distribution [38]. Thus, we first calibrate PLM on a given prompt and then use calibrated class scores for classification.
Specifically, we use a template to wrap empty input $x_{e}=$“” and query the model with $\mathcal{T}_{j}(x_{e})$ to obtain the predictions of label words $\mathbf{s}_{ej}\in\mathbb{R}^{N|\emph{S}|}$ to calibrate. Then, given a prediction $\mathbf{s}_{ij}$ for sample $x_{i}$ wrapped by prompt $\mathcal{T}_{j}$, we can normalize them by $\mathbf{s}_{ej}/\text{mean}(\mathbf{s}_{ej})$. The calibration class scores are calculated as follows:
$\mathbf{\bar{s}}_{ij}=\text{diag}(\mathbf{s}_{ej}/\text{mean}(\mathbf{s}_{ej}))^{-1}\mathbf{s}_{ij}$ (8)
where $\mathbf{\bar{s}}_{ij}\in\mathbb{R}^{N|\emph{S}|}$ is calibrated prediction over the expanded label set. After calibration, PLMs predict a uniform distribution for context-independent inputs such as empty input on expanded label words.
Then, we use the weight vector on the expanded label set S to aggregate the predictions to form the calibrated class scores $\mathbf{\hat{s}}_{ij}\in\mathbb{R}^{N}$. Finally, we average the class scores of multiple prompts to get averaged calibrated class scores $\mathbf{\hat{s}}_{i}$ for classification.
$\mathbf{\hat{s}}_{i}=\frac{\sum_{j=1}^{P}\mathbf{\hat{s}}_{ij}}{P}$ (9)
IV-C3 Joint Decoding
Finally, we sum OT scores and averaged calibrated class scores to jointly decode the predictions.
$\hat{y}_{i}=\hat{y}^{OT}_{i}+\beta\mathbf{\hat{s}}_{i}$ (10)
where $\hat{y}_{i}$ is the final predictions, $\hat{y}^{OT}_{i}$ = [Dec($x_{i}$,1), $\cdots$, Dec($x_{i}$,$N$)] $\in\mathbb{R}^{N}$ denotes OT scores, and $\beta$ is a hyper-parameter to balance model and prior knowledge inside PLM.
IV-D Summary of MPD
In summary, there are three steps to train our model. In the first step, we use multiple prompts and multiple class prototypes to obtain text representations and class representations. In the second step, we use optimal transport to align text representations with prototypes and calculate the optimal transport distance. In the final step, we use cross-entropy loss with OT distances as logits to optimize parameters. The whole training flow is illustrated in Figure 2 and Algorithm 1.
During the inference phase, we first expand label words and get calibrated class scores. Then, we sum OT scores and calibrated class scores for classification.
V Experiments
To evaluate the proposed MPD, we conduct extensive experiments to answer the following Research Questions (RQ):
  • RQ1: Whether our proposed model outperforms existing MaaS methods?
  • RQ2: How does the number of trainable parameters and training time of the model compare to the baselines?
  • RQ3: How does each of the components of MPD contribute to the final performance?
  • RQ4: How do the number and type of prompts affect the performance?
  • RQ5: Can our method be transferred to other PLM architectures?
Thereafter, we also explore the effects of hyper-parameters, the effectiveness of OT, and visualize the expanded label words.
V-A Datasets
We conduct experiments on several common natural language understanding tasks including sentiment analysis, topic classification, and natural language inference (NLI). The statistics of tasks and datasets are shown in Table II.
The sentiment analysis task includes five datasets. The Stanford Sentiment Treebank (SST2) dataset [39], IMDB dataset [40], and MR [41] were both extracted from movie reviews. The Yelp dataset [42] consists of reviews from Yelp. The CR dataset [43] was extracted from customer reviews.
The topic classification task includes three datasets. AG’s news dataset [42] collects more than 1 million news articles from more than 2,000 news sources. We also only use the title and description fields to classify news. The Yahoo dataset [42] was collected from the Yahoo Webscope program and contains questions and their answers. The fields we used include question title, question content, and best answer. The DBPedia ontology classification dataset [42] is constructed by picking 14 non-overlapping classes from DBPedia 2014 derived from Wikipedia.
The NLI task includes three datasets. The Recognizing Textual Entailment (RTE) dataset [44] comes from a series of textual entailment challenges. Examples on the RTE dataset are constructed based on news and Wikipedia text. The Stanford Natural Language Inference (SNLI) corpus [45] is a collection of human-written English sentence pairs manually labeled for balanced classification with the labels entailment, contradiction, and neutral. MNLI (Multi-genre Natural Language Inference) [46] offers ten distinct genres of written and spoken English NLI data. MNLI-m and MNLI-mm correspond to versions matched and mismatched of the MNLI dataset, respectively.
We also randomly sample $N$ = 1, 4, 16 samples for each class from the original training set to construct the few-shot training set.
TABLE II: Statistics of datasets.
Task Dataset # Class # Test
Sentiment Analysis SST2 2 872
Yelp 2 38,000
IMDB 2 25,000
MR 2 2001
CR 2 2000
Topic Classification AG’s News 4 7,600
Yahoo 10 60,000
DBPedia 14 70,000
NLI RTE 2 277
SNLI 3 9,842
MNLI-m 3 9,815
MNLI-mm 3 9,832
TABLE III: Experiment results for MaaS adaptation methods. The results (i.e., average accuracy and standard deviation (%)) over 5 runs. The best results are in bold.
$N$ Method SST2 IMDB Yelp AG DB Yahoo RTE SNLI MNLI-m/mm Avg.
1 ICL $81.5_{3.7}$ $65.6_{11.4}$ $81.1_{10.6}$ $66.7_{4.8}$ $71.7_{2.6}$ $53.2_{6.2}$ $45.0_{4.7}$ $46.1_{5.3}$ ${\bf 53.6_{0.5}}/53.9_{0.8}$ $61.8_{5.1}$
RLPrompt $63.5_{6.3}$ $65.0_{6.5}$ $66.3_{6.9}$ $72.5_{4.5}$ $65.6_{5.5}$ $38.1_{5.8}$ $53.8_{5.3}$ $36.5_{3.0}$ $40.3_{2.0}/41.0_{2.1}$ $54.3_{4.8}$
BBT $83.4_{1.3}$ $89.0_{0.1}$ $89.7_{0.1}$ $75.4_{0.8}$ $59.1_{1.7}$ $31.2_{2.7}$ $52.3_{1.4}$ $38.5_{0.8}$ $43.4_{2.5}/42.9_{3.3}$ $60.5_{1.5}$
BBTv2 $83.3_{2.5}$ $89.0_{0.2}$ $89.9_{0.2}$ $74.3_{3.2}$ $74.2_{5.2}$ $34.0_{3.5}$ $48.2_{5.7}$ $38.6_{4.0}$ $44.2_{3.2}/44.3_{4.5}$ $62.0_{3.2}$
PromptBoosting $86.7_{2.6}$ $82.4_{6.1}$ $88.7_{2.5}$ $58.7_{11.8}$ $73.0_{4.8}$ $23.7_{7.0}$ $50.0_{5.9}$ $43.5_{6.1}$ $36.8_{1.6}/36.3_{2.3}$ $58.0_{5.1}$
DecT $90.8_{0.2}$ $91.2_{0.3}$ $94.8_{0.1}$ $79.9_{1.1}$ $78.8_{0.9}$ ${\bf 55.2_{0.8}}$ $56.0_{2.7}$ ${\bf 47.7_{4.1}}$ $52.2_{2.7}/53.3_{3.0}$ $70.0_{1.6}$
MPD ${\bf 92.3_{0}}$ ${\bf 94.4_{0.3}}$ ${\bf 95.6_{1.1}}$ ${\bf 83.2_{1.5}}$ ${\bf 84.4_{3.5}}$ ${53.6_{0.9}}$ ${\bf 57.6_{1.7}}$ ${46.6_{4.1}}$ ${53.1_{1.5}/\bf 54.1_{1.8}}$ ${\bf 71.5_{1.6}}$
4 ICL $60.3_{9.8}$ $80.4_{6.6}$ $77.4_{14.6}$ $65.1_{5.4}$ $71.7_{6.5}$ $49.9_{9.9}$ $42.7_{3.9}$ $42.1_{3.2}$ $44.7_{5.9}/45.2_{6.0}$ $58.0_{7.2}$
RLPrompt $80.7_{7.5}$ $75.8_{10.1}$ $78.8_{7.3}$ $76.1_{4.8}$ $76.3_{5.9}$ $45.0_{3.1}$ $53.5_{2.9}$ $36.3_{2.6}$ $44.4_{2.9}/45.5_{3.8}$ $61.2_{5.1}$
BBT $84.5_{1.2}$ $89.8_{0.9}$ $90.2_{0.6}$ $79.0_{2.1}$ $67.7_{3.5}$ $42.9_{0.6}$ $48.4_{4.0}$ $40.5_{1.3}$ $41.2_{1.7}/40.7_{2.0}$ $62.5_{1.8}$
BBTv2 $86.6_{2.2}$ $89.4_{0.6}$ $90.3_{0.5}$ $79.1_{2.1}$ $89.0_{1.7}$ $46.0_{1.4}$ $46.2_{2.3}$ $40.8_{4.3}$ $44.0_{0.9}/44.8_{1.6}$ $65.6_{1.8}$
PromptBoosting $88.9_{2.3}$ $83.0_{5.2}$ $92.3_{2.1}$ $78.2_{6.8}$ $90.1_{0.7}$ $36.4_{5.1}$ $53.5_{5.9}$ $53.4_{3.4}$ $39.8_{4.5}/40.3_{5.7}$ $65.6_{4.2}$
DecT $87.6_{1.6}$ $89.6_{0.9}$ $94.8_{0.7}$ $81.9_{2.6}$ $89.1_{0.6}$ $59.9_{2.1}$ $56.7_{2.7}$ $53.2_{2.9}$ $52.2_{2.3}/53.4_{2.4}$ $71.8_{1.9}$
MPD ${\bf 92.6_{0.1}}$ ${\bf 94.9_{0.2}}$ ${\bf 96.3_{0.3}}$ ${\bf 85.9_{1.0}}$ ${\bf 92.8_{3.0}}$ ${\bf 62.2_{1.3}}$ ${\bf 59.2_{3.6}}$ ${\bf 57.1_{2.4}}$ ${\bf 54.4_{0.1}/56.5_{0.6}}$ ${\bf 75.2_{1.1}}$
16 ICL $71.5_{15.8}$ $80.6_{6.0}$ $73.7_{14.5}$ $64.4_{6.0}$ $71.8_{9.1}$ $52.6_{5.7}$ $43.8_{7.0}$ $42.0_{6.3}$ $51.4_{3.0}/52.1_{3.3}$ $60.0_{7.7}$
RLPrompt $87.0_{2.6}$ $87.6_{2.4}$ $95.1_{1.0}$ $80.2_{0.7}$ $80.8_{3.3}$ $48.1_{2.2}$ $54.3_{2.8}$ $41.1_{5.0}$ $43.3_{3.9}/44.3_{4.5}$ $66.2_{2.8}$
BBT $89.6_{0.3}$ $89.3_{0.4}$ $91.5_{0.2}$ $81.5_{0.8}$ $87.8_{3.0}$ $48.3_{1.4}$ $52.6_{2.2}$ $46.6_{1.3}$ $40.0_{2.6}/39.9_{2.9}$ $66.7_{1.5}$
BBTv2 $90.3_{1.7}$ $88.6_{2.1}$ $92.9_{0.6}$ $85.3_{0.5}$ $93.6_{0.7}$ $52.0_{1.4}$ $56.7_{3.3}$ $57.3_{2.3}$ $50.1_{2.4}/51.7_{3.2}$ $71.9_{1.8}$
PromptBoosting $87.6_{3.0}$ $86.2_{3.1}$ $94.7_{1.0}$ $85.2_{0.9}$ $95.0_{0.5}$ $46.6_{2.4}$ $60.0_{5.5}$ $61.3_{3.5}$ $52.5_{1.5}/50.4_{5.1}$ $72.0_{2.7}$
DecT $91.0_{0.5}$ $91.0_{0.9}$ $95.4_{0.3}$ $86.4_{0.4}$ $94.6_{0.5}$ $64.2_{0.7}$ $59.7_{1.8}$ $60.5_{0.8}$ $55.3_{1.3}/56.8_{1.5}$ $75.5_{0.9}$
MPD ${\bf 91.9_{0.1}}$ ${\bf 95.1_{0.1}}$ ${\bf 96.4_{0.1}}$ ${\bf 87.9_{0.4}}$ ${\bf 96.7_{0.2}}$ ${\bf 68.3_{0.3}}$ ${\bf 61.7_{2.8}}$ ${\bf 62.4_{1.3}}$ ${\bf 57.5_{1.1}/59.7_{1.1}}$ ${\bf 77.8_{0.8}}$
V-B Baselines
We compare with some strong MaaS methods.
  • In-context learning (ICL) [47] concatenates some text-label pairs before the test samples.
  • RLPrompt [9] optimizes discrete prompts with RL and adopts soft Q-learning to find the optimal prompt.
  • BBT [3] optimizes soft prompt tokens with a derivative-free evolutionary algorithm.
  • BBTv2 [8] further prepends continuous prompts to every layer and uses a divide-and-conquer algorithm to optimize them.
  • PromptBoosting [10] sequentially learns multiple weak learners and adopts AdaBoost to ensemble them.
  • DecT [11] extracts text feature and optimizes class prototypes as hyperspheres with an additional radius parameter.
  • TEMPERA [27] designs the agent to perform different editing techniques to construct query-dependent prompts.
  • DP2O [28] first designs a prompt generation strategy based on GPT-4 and then designs a reinforcement learning framework based on policy gradients to match suitable prompts for a single input.
V-C Implementation Details
[…]
TABLE IV: Comparsion with TEMPERA and DP2O following their experimental settings under the 16-shot setting. The results (i.e., average accuracy and standard deviation (%)) over 4 runs. The best results are in bold.
Method SST2 MR CR Yelp Avg.
TEMPERA $91.9_{2.0}$ $88.0_{1.1}$ $91.1_{1.6}$ $92.6_{1.7}$ $90.9_{1.6}$
$\bf{DP_{2}O}$ $\bf{93.6_{0.7}}$ $88.6_{0.9}$ $90.8_{0.5}$ $94.3_{0.4}$ $91.8_{0.6}$
MPD $92.1_{0.1}$ $\bf{89.6_{0.2}}$ $\bf{91.3_{0.3}}$ $\bf{96.4_{0.1}}$ $\bf{92.4_{0.2}}$
TABLE V: Efficiency comparison for MaaS methods under 16-shot setting on two datasets.
Method Trainable Param (K) Query Number SST2 AG
Acc Training Time (s) Acc Training Time (s)
ICL 0 0 71.5 0 64.4 0
RLPrompt 3100 12,000 87.0 29012 80.2 97781
BBT 0.5 8,000 89.6 1619 87.8 5228
BBTv2 12 8,000 90.3 1624 85.3 6546
PromptBoosting 0.4 10 87.6 190 85.2 398
DecT 130 1 91.0 1.4 86.4 2.3
Ours 132 3 91.9 3.5 87.9 6.7
Method Trainable Param (K) Query Number SST2 MR
Acc Training Time (s) Acc Training Time (s)
TEMPERA 16898 7800 91.9 69418 88.0 71178
DP2O 609 15 93.6 159 88.6 159
Ours 132 3 92.1 3.5 89.6 3.6
V-D Results (RQ1)
Table III shows the performance on 10 datasets under the few-shot setting following Cui et al. [11]. We can see that our model achieves state-of-the-art performance. Specifically, compared to DecT, our approach achieves 1.5%, 3.4%, and 2.3% average performance improvement under 1, 4, 16-shot settings. Among all settings on all datasets, our model performs the best in 27 out of the 30. This shows our model is effective and the improvement is consistent. It is worth noting that our model performs well on both simple (i.e., sentiment analysis and topic classification) and difficult tasks that require semantic understanding (i.e., NLI). This shows that the improvement is task-independent.
By observing other baseline methods, ICL is a strong baseline under the 1-shot setting. Due to the length limitation of PLM, ICL does not achieve performance improvement with increasing shot number. BBT and RLPrompt achieve better performance compared to ICL under the 4-shot and 16-shot settings. This illustrates the effectiveness of learning prompts when a sufficient number of training sets are available. BBTv2 exceeds BBT and RLPrompt, indicating the effectiveness of inserting continuous prompts into every layer of the PLM. PromptBoosting and DecT usually achieve better performance than BBT and BBTv2. This is because these two methods do not need to search prompts in a large space and can benefit from gradients.
In addition, our method also achieves the minimum average standard deviation across different datasets and settings. This shows that our approach is the most robust. As the shot number increases, our model achieves the minimum standard deviation on 1, 6, and 8 datasets. This shows that compared to other methods, our model is more likely to obtain a stable result as the shot number increases.
We further compare MPD with two baselines (i.e., TEMPERA and DP2O) following their experimental setting on four datasets in Table IV. Our results demonstrate that our model consistently outperforms these baselines in terms of both average performance and standard deviations. Specifically, compared to TEMPERA and DP2O, our approach achieves 1.5% and 0.6% average performance improvement. This illustrates that using multiple prompts for joint learning is more efficient and robust. Both TEMPERA and DP2O belong to input-side adaptation and achieve strong performance. This shows that constructing query-dependent prompts and matching prompts are effective. It is worth noting that the inconsistency between the results of SST2 in Table III and Table IV is due to the use of different experimental settings in the two tables.
TABLE VI: Experiment results for more training data over 5 runs. Fine-tuning method requires gradients of the model to update the model.
$N$ Method SST2 IMDB Yelp AG DB Yahoo RTE SNLI MNLI-m/mm Avg.
64 Fine-tuning $92.5_{1.9}$ $86.3_{3.8}$ $94.5_{1.4}$ $87.4_{0.6}$ $98.2_{0.2}$ $69.0_{0.7}$ $67.7_{3.2}$ $66.6_{6.4}$ $65.6_{2.9}/67.7_{4.0}$ $79.6_{2.5}$
DecT $92.4_{0.5}$ $91.3_{0.5}$ $94.9_{0.5}$ $89.2_{0.3}$ $97.0_{0.1}$ $69.3_{0.4}$ $65.7_{1.7}$ $67.2_{1.0}$ $62.0_{1.4}/63.3_{1.3}$ $79.2_{0.8}$
MPD $92.0_{0.2}$ $95.1_{0.5}$ $96.6_{0.1}$ $89.6_{0.3}$ $97.4_{0.1}$ $70.6_{0.3}$ $67.6_{1.2}$ $68.2_{0.7}$ $63.4_{0.7}/65.1_{1.1}$ $80.6_{0.5}$
256 Fine-tuning $92.0_{0.9}$ $92.1_{0.2}$ $94.3_{0.3}$ $89.6_{0.3}$ $98.5_{0.2}$ $70.2_{0.4}$ $79.8_{1.0}$ $84.4_{0.4}$ $77.2_{0.2}/78.7_{0.3}$ $85.7_{0.4}$
DecT $92.7_{0.2}$ $92.1_{0.1}$ $95.6_{0.1}$ $90.3_{0.1}$ $97.4_{0.1}$ $71.3_{0.1}$ $69.2_{1.0}$ $69.7_{0.4}$ $68.0_{0.3}/69.4_{0.3}$ $81.6_{0.3}$
MPD $92.3_{0.2}$ $95.1_{0.1}$ $96.7_{0}$ $90.6_{0.2}$ $97.6_{0.1}$ $71.4_{0.2}$ $71.6_{1.0}$ $70.3_{0.6}$ $68.9_{0.4}/69.9_{0.7}$ $82.4_{0.4}$
V-E Efficiency Comparison (RQ2)
We further compare our model with baselines in terms of efficiency, including the number of trainable parameters, the number of queries to access PLM, and training time.
As shown in Table V, our model is effective and efficient. The efficiency of our approach significantly outperforms all baselines except DecT. Specifically, our model queries PLM only three times per training data, which is about 3$\times$, 5$\times$, 2600$\times$, and 2700$\times$ less than PromptBoosting, DP2O, TEMPERA, and BBTv2, respectively. In addition, the training time of our method is about 45$\times$ and 50$\times$ faster than that of DP2O and PromptBoosting. This further illustrates that the trained parameters in our method are lightweight. Our method uses multiple prompts, so the training time is slightly longer than in DecT. Overall, the training speed of our method is still fast. In addition, our model only includes 132K trainable parameters, significantly smaller than some baselines (i.e., RLPrompt, TEMPERA, and DP2O). Even though some methods (i.e., BBT, BBTv2, and PromptBoosting) have fewer trainable parameters, our model takes less time to train and achieves better performance.
V-F Performance on More Training Data
We also explore the scalability of our model beyond the few-shot setting, we conduct experiments with a larger training set (i.e., $N=$ 64 and 256). We compare our model with DecT and a strong baseline fine-tuning. Fine-tuning requires gradients of PLM to update full parameters of PLM and is a strong baseline method under the MaaS setting.
As shown in Table VI, firstly, our model and DecT both achieve performance improvement when the shot number increases. This shows that these two methods can still fit larger datasets even with a small number of tunable parameters. In addition, as the shot number increases, the average standard deviation of the two models decreases. This suggests that more data helps to learn a robust model. Secondly, our model also exceeds DecT when $N$ is large. This further shows that our model is effective and the improvement is consistent. It is worth noting that the gap between the two methods is narrowing as the shot number increases. This may be due to the similar number of trainable parameters for both models. Thirdly, our method exceeds fine-tuning in the 64-shot setting and is exceeded by fine-tuning in the 256-shot setting. This is because our method cannot update PLM parameters, it can only update a small portion of parameters. Thus, when the shot number is larger, our model may not be able to learn more knowledge. Finally, similar to DecT, our model does not show significant performance differences compared to fine-tuning in simple tasks (i.e., sentiment analysis and topic classification). However, there are significant differences in harder tasks (i.e. NLI). This suggests that harder tasks require more parameters to learn corresponding semantics.
Figure 4: Effects of prompts on AG’s news dataset under 16-shot setting. 1, 2, 3, 4, and 5 represent 5 different prompts. 1+2 denotes we use two prompts (i.e., 1 and 2) to train the model. The results over 5 runs.
TABLE VII: The ablation study of our model on 10 datasets in Table III. We run each experiment over 5 random seeds.
Model Setting Average Accuracy
1 4 16
Ours 71.5 75.2 77.8
w/o Class Scores 57.8 69.6 76.5
w/o OT Scores 68.0 68.0 68.0
w/o Expansion 70.7 74.8 77.7
V-G Ablation Study (RQ3)
We conduct ablation study to validate the effectiveness of each component shown in Table VII.
Firstly, all components achieve performance improvement. This shows all proposed components are effective. Secondly, the performance of w/o Class Scores (i.e., only using OT scores) improves rapidly as the shot number increases. This shows that shot number has a significant effect on OT scores. Learned OT scores are more effective when the shot number is large. Thirdly, w/o OT Scores (i.e., only averaging calibrated class scores of multiple prompts) also achieves good performance, especially in the 1-shot setting. This shows that averaging the class scores of multiple prompts is also a strong baseline, especially in the 1-shot setting. Finally, label words expansion achieves more significant performance gains when the shot number is small. This is because label words expansion improves the performance of class scores and the model relies more on class scores when the shot number is small.
TABLE VIII: Five templates used in our experiments. The fourth prompt comes from Deng et al. [9].
Template Id Template
1 [ Topic : [MASK] ] $x_{1}$ $x_{2}$
2 [ Category : [MASK] ] $x_{1}$ $x_{2}$
3 $x_{1}$ $x_{2}$ The topic is about [MASK].
4 [MASK] Alert Blog Dialogue Diary Accountability $x_{1}$ $x_{2}$
5 A [MASK] news : $x_{1}$ $x_{2}$
V-H Effects of Prompts (RQ4)
We explore the effects of prompts by changing the combination of prompts in Figure 4. Specifically, we use different prompts and change the number of prompts. The five templates are shown in Table VIII.
Firstly, as the prompt number increases, performance will be better and more stable. When the number of prompts exceeds 3, the improvement is marginal. This shows that our model is effective and robust. Secondly, changing specific prompts has less performance fluctuation with multiple prompts compared to using a single prompt. Specifically, when we use one/two/three/four prompts, the change in performance is 1.7%/0.3%/0.2%/0.2%. This shows that our method is robust across multiple prompts. Thirdly, the good performance of a prompt’s class scores does not mean that using it can achieve better performance. This shows that there is a gap between the representation of hidden states and class scores.
TABLE IX: Three different types of templates (i.e., vanilla prompt, instruction prompt, and example prompt) used in our experiments.
Template Type Template
V [ Topic : [MASK] ] $x_{1}$ $x_{2}$
I Classify the news articles into the categories of Politics, Sports, Business, and Technology. $x_{1}$ $x_{2}$ This topic is about [MASK].
E Article: Venezuelans Vote Early in Referendum on Chavez Rule (Reuters). Reuters - Venezuelans turned out early and in large numbers on Sunday to vote in a historic referendum that will either remove left-wing President Hugo Chavez from office or give him a new mandate to govern for the next two years. Answer: Politics. Article: $x_{1}$ $x_{2}$ Answer: [MASK].
TABLE X: Performance (16-shot) for our method on different scales of T5 [50] and LLaMA2 [51]. We also run each experiment over 5 random seeds and report the average accuracy and standard deviation (%).
Model SST2 IMDB AG DB RTE SNLI Avg.
[…]
TABLE XI: Accuracy using various types of prompts on AG’s news dataset under the 16-shot setting.
        Adopted Prompts         Acc
        V         87.44
        I         86.06
        E         87.47
        V+I         87.49
        V+E         88.11
        I+E         87.66
        V+I+E         87.92
V-I Effectiveness on Various Types of Prompts (RQ4)
In this section, we further show the effectiveness on various types of prompts, including instruction and example. Specifically, we use the vanilla prompt 1 in Table VIII (denoted as V), instruction (denoted as I), and example (denoted as E, we use the first training data from the AG’s news dataset to construct example prompt) on the AG’s news dataset under the 16-shot setting. Three prompts as shown in Table IX.
Firstly, our method usually achieves performance improvement compared to using a single prompt in Table XI. Specifically, using two prompts exceeds using either prompt individually. For example, V+I exceeds V and I, respectively. It is worth noting that V+E exceeds all results in Figure 4. This shows that using two different types of prompts sometimes yields better results. Secondly, the performance fluctuation of using two prompts is smaller than that of a single prompt. Thirdly, V+I+E does not exceed V+I due to the poor performance of I. This is because RoBERTa-large has not been fine-tuned by instruction-tuning.
V-J Extension to Other PLMs (RQ5)
[…]
TABLE XII: The results of label words expansion.
Dataset Label Words Label Words Expansion
SST2, IMDB Yelp bad ’bad’, ’Bad’, ’ bad’, ’ Bad’, ’good’, ’poor’, ’wrong’, ’big’, ’worst’, ’ BAD’
great ’great’, ’Great’, ’ great’, ’small’, ’huge’, ’little’, ’good’, ’perfect’, ’powerful’, ’big’
AG’s News politics ‘politics’, ‘Politics’, ‘political’, ‘Political’, ‘policy’, ‘ politics’, ‘ Politics’, ‘democracy’, ‘history’, ‘business’
technology ‘technology’,‘Technology’, ‘tech’, ‘software’, ‘science’, ‘ technology’, ‘computer’, ‘engineering’, ‘technical’, ‘device’
business ‘business’, ‘Business’, ‘ business’, ‘ Business’, ‘commercial’, ‘property’, ‘biz’, ‘office’, ‘trade’, ‘everything’
sports ‘sports’, ‘Sports’, ‘Sport’, ‘football’, ‘ sports’, ‘ Sports’, ‘music’, ‘gaming’, ‘games’, ‘military’
Yahoo soc ’soc’, ’Soc’, ’social’, ’community’, ’offic’, ’civil’, ’ soc’, ’sn’, ’mom’, ’ Soc’
science ’science’, ’Science’, ’scientific’, ’ science’, ’technology’, ’biology’, ’research’, ’history’, ’scient’, ’evidence’
health ’health’, ’Health’, ’ health’, ’medical’, ’healthy’, ’ Health’, ’hospital’, ’Medical’, ’cancer’, ’safety’
education ’education’, ’Education’, ’ education’, ’learning’, ’training’, ’educated’, ’awareness’, ’student’, ’college’, ’school’
com ’com’, ’Com’, ’ com’, ’COM’, ’ Com’, ’org’, ’comm’, ’co’, ’net’, ’gov’
sports ’sports’, ’Sports’, ’Sport’, ’football’, ’ sports’, ’ Sports’, ’music’, ’gaming’, ’games’, ’military’
business ’business’, ’Business’, ’ business’, ’ Business’, ’commercial’, ’property’, ’biz’, ’office’, ’trade’, ’everything’
ent ’ent’, ’ENT’, ’ents’, ’enting’, ’ented’, ’ant’, ’ente’, ’ento’, ’ency’, ’enta’
family ’family’, ’Family’, ’ family’, ’community’, ’ Family’, ’daughter’, ’brother’, ’parents’, ’mom’, ’small’
politics ’politics’, ’Politics’, ’political’, ’Political’, ’policy’, ’ politics’, ’ Politics’, ’democracy’, ’history’, ’business’
DBPedia company ’company’, ’Company’, ’ company’, ’project’, ’community’, ’family’, ’Companies’, ’city’, ’business’, ’office’
school ’school’, ’School’, ’ school’, ’college’, ’ School’, ’chool’, ’fashioned’, ’education’, ’student’, ’program’
artist ’artist’, ’Artist’, ’ artist’, ’ Artist’, ’music’, ’creator’, ’director’, ’manager’, ’editor’, ’student’
ath ’ath’, ’aths’, ’athe’, ’ATH’, ’athi’, ’atha’, ’athy’, ’oth’, ’athing’, ’athan’
politics ’politics’, ’Politics’, ’political’, ’Political’, ’policy’, ’ politics’, ’ Politics’, ’democracy’, ’history’, ’business’
trans ’trans’, ’Trans’, ’ Trans’, ’ trans’, ’rans’, ’poly’, ’simple’, ’project’, ’small’, ’digital’
building ’building’, ’build’, ’builder’, ’Building’, ’builders’, ’ building’, ’built’, ’ Building’, ’making’, ’fighting’
river ’river’, ’River’, ’rider’, ’ river’, ’roller’, ’lake’, ’city’, ’later’, ’country’, ’current’
vill ’vill’, ’Vill’, ’ Vill’, ’vell’, ’ville’, ’vag’, ’font’, ’coll’, ’christ’, ’ vill’
animal ’animal’, ’Animal’, ’ animal’, ’species’, ’human’, ’monster’, ’ Animal’, ’horse’, ’baby’, ’adult’
plant ’plant’, ’ plant’, ’ Plant’, ’bomb’, ’ Plants’, ’flower’, ’forest’, ’train’, ’fruit’, ’print’
album ’album’, ’ album’, ’ Album’, ’music’, ’movie’, ’episode’, ’concert’, ’download’, ’artist’, ’film’
film ’film’, ’Film’, ’movie’, ’ film’, ’ Film’, ’Movie’, ’ Films’, ’music’, ’video’, ’picture’
book ’book’, ’books’, ’Book’, ’ book’, ’ Book’, ’BOOK’, ’Books’, ’ books’, ’sheet’, ’ebook’
Figure 5: Effects of $\beta$ over 5 runs on two datasets.
V-K Effects of Hyper-Parameter $\beta$
In this section, we explore the effects of $\beta$ on two datasets in Figure 5. Hyper-parameter $\beta$ controls the ratio of OT scores and class scores.
Firstly, our model needs a larger $\beta$ under the 1-shot setting. This is because the amount of data is small and the model relies heavily on class scores for prediction rather than OT scores. Thus, a larger $\beta$ usually achieves better performance under the 1-shot setting. Secondly, as the shot number increases, the effect of $\beta$ on the results diminishes. This is because as the shot number increases, the performance of OT scores is enhanced and relies more on the OT scores for prediction. Thirdly, as the shot number increases, the standard deviation becomes smaller. This shows that the OT scores are more stable when the training size increases.
Figure 6: Effects of $Q$ over 5 runs on two datasets.
V-L Effects of the Number of Prototypes
In this section, we explore the effects of the number of prototypes on two datasets in Figure 6.
Firstly, we can see that using multiple prototypes tends to outperform using one. This shows that using multiple prototypes is effective in our multiple prompts setting. Secondly, the improvement is more significant under the 4-shot and 16-shot. This shows that as the shot number increases, multiple prototypes are more effective. Finally, as the shot number increases, a large $Q$ usually performs better. For example, the model achieves the best performance on two datasets under the 4-shot setting with $Q=4$ and 3. The model achieves the best performance on two datasets under the 4-shot setting with $Q=5$ and 4. This is because when the shot number increases, the model needs a large $Q$ to capture subtle differences between multiple texts of the same class. We simply set $Q=3$ on all datasets and settings to avoid hyper-parameter selection.
TABLE XIII: Performance of two variants to replace optimal transport plan. We run each experiment over 5 random seeds.
Model Setting Average Accuracy
1 4 16
Ours 71.5 75.2 77.8
Uniform Distribution 70.7 74.7 77.1
Cosine Similarity 68.3 73.8 77.4
V-M Visualization of Label Words Expansion
We show the expanded label words in Table XII. We can see that our proposed simple expansion method expands some similar words and some semantically related words. For example, on the three sentiment datasets (i.e., SST2, IMDB, and Yelp), label word ‘bad’ expands some semantically relevant words (i.e., ‘poor’, ‘wrong’, and ‘worst’), and label word ‘great’ expands some relevant words (i.e., ‘good’, ‘perfect’, and ‘powerful’). On the AG’s News dataset, label word ‘politics’ expands ‘democracy’ and ‘history’, and label word ‘technology’ expands ‘software’, ‘science’, and so on. This shows that our proposed label words expansion is effective. While ‘bad’ expands ‘good’, we assign a weight to ‘good’ based on its cosine similarity to the target class word to avoid the effect of noise.
V-N Effectiveness of OT
We use two variants to illustrate the effectiveness of fine-grained matching using OT, including uniform distribution and cosine similarity.
We can see that OT achieves the best performance in Table XIII. This shows that OT can effectively match the relationship between text representations and prototypes. The uniform distribution does not take into account the similarity between representations and prototypes. Cosine similarity only calculates the similarity between two elements and cannot take into account the similarity with other elements.
VI Conclusion
In this paper, we propose the multi-prompting decoder framework for output-side MaaS adaptation on few-shot tasks. To achieve better decoding performance, we design two decoding strategies to decode PLMs’ output hidden states and class scores, respectively. Extensive experiments demonstrate that our method achieves new state-of-the-art results under the few-shot setting on multiple natural language understanding tasks, including sentiment analysis, topic classification, and natural language inference. Besides, our experiments also show that our model is lightweight, efficient, and applicable to PLMs of different architectures (e.g., encoder only, decoder only, and encoder-decoder).